### v2

# Artificial Intelligence LLM Large Language Model

Abstract.
As the use of Large Language Model (LLM) chatbots by students and researchers becomes more prevalent, universities are pressed to develop AI strategies. One strategy that many universities pursue is to customize pre-trained LLM as-a-service (LLMaaS). While most studies on LLMaaS chatbots prioritize technical adaptations, we focus on psychological effects of user-salient customizations, such as interface changes. We assume that such customizations influence users’ perception of the system and are therefore important in guiding safe and appropriate use. In a field study, we examine how students and employees ($N=526$) at a German university perceive and use their institution’s customized LLMaaS chatbot compared to ChatGPT. Participants using both systems ($n=116$) reported greater trust, higher perceived privacy and less experienced hallucinations with their university’s customized LLMaaS chatbot in contrast to ChatGPT. We discuss theoretical implications for research on calibrated trust, and offer guidance on the design and deployment of LLMaaS chatbots.
Large Language Models as-a-Service (LLMaaS), Customization, Trust, Hallucinations, University
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Human-centered computing Natural language interfacesccs: Security and privacy Privacy protections
1. Introduction
Large Language Model (LLM) chatbots are becoming increasingly prevalent in educational settings. Current research shows high adoption rates of commercial off-the-shelf LLM chatbots (e. g., ChatGPT) among students and university staff (Zhang et al., 2024; Zhou et al., 2023; von Garrel and Mayer, 2023). Consequently, universities face mounting pressure to develop strategies for AI integration (Kapania et al., 2025). This involves multiple decisions, such as whether and how to provide employees with access to LLMs. To guide the informed and secure use of such systems, many organizations, including universities, corporations, and government agencies, are exploring the deployment of LLM as-a-Service (LLMaaS) as a feasible and cost-effective way to adopt AI in-house (La Malfa et al., 2023; Chkirbene et al., 2024). These services can form the basis for internal LLMaaS-based chatbots in organizations. The adoption of such services is driven by several practical considerations, including the opportunity to offer access to all university members while maintaining institutional oversight and alignment with organizational policies (Paavola, 2024). One of the main benefits of AI-as-a-Service (AIaaS) solutions, of which LLMaaS is a specific instance, is the flexibility they offer in customizing various aspects (Diaferia et al., 2022), such as fine-tuning models or displaying user interfaces in corporate design.
Customizations of LLMaaS solutions play a critical role, as they can enhance model performance within specific organizational contexts and strengthen security and privacy measures (Sundberg and Holmström, 2023; Lewicki et al., 2023). However, even choices that do not impact model performance or security, such as branding the user interface, could lead to significant changes in how the system is perceived and used. Psychological research suggests that people draw inferences about a system’s trustworthiness based on indicators available to them, also referred to as cues (de Visser et al., 2014), a mechanism formalized in the Trustworthiness Assessment Model (TrAM) (Schlicker et al., 2025). Real-world incidents show the importance of understanding user perception of LLM chatbots: cases like the data leak by Samsung employees via ChatGPT illustrate how overly trusting a system or having false privacy expectations can lead to the unintended disclosure of sensitive information (Mauran, 2023). Such incidents underscore the role of trust calibration, meaning the degree to which users’ perceived trustworthiness of a system aligns with its actual properties and limitations (Lee and See, 2004; Parasuraman and Riley, 1997). While recent studies have focused on how to intentionally improve calibrated trust and reliance on LLMs through interventions (Bo et al., 2025; Kim et al., 2024; Ma et al., 2025a; Vasconcelos et al., 2025), we lack research on how customization choices of real-world LLM chatbots impact the calibration of user perceptions. In particular, research is lacking on customization choices that are salient to users (e. g., in the form of cues perceivable in the user interface) but have no impact on aspects such as the model, training data, privacy safeguards, or system infrastructure, which have been the primary focus of prior work (Diaferia et al., 2022; Ali et al., 2019; Tönjes et al., 2024).
In this work, we examine if cues – emerging from different customization choices in LLMaaS chatbots – influence users’ perceptions and usage. We draw on the concept of trust calibration (Lee and See, 2004; Parasuraman and Riley, 1997) and the TrAM (Schlicker et al., 2025) to empirically investigate the alignment (or misalignment) between users’ perception of a customized LLMaaS and the system’s actual characteristics, thereby answering the following research question:
RQ: How do customizations of LLMaaS chatbots impact end users’ trust, perception, and usage of LLM chatbots in professional and academic contexts compared to commercial AI alternatives?
To examine this research question, we conducted a survey at a major German university. A diverse sample of students and university employees (N = 526) answered questions about their experiences with the university’s customized LLMaaS chatbot and with ChatGPT (the most widely used LLM chatbot among German students (von Garrel and Mayer, 2025)). We measured trust, privacy concerns, cautious behavior toward LLM hallucinations, experienced LLM hallucinations, and sustainability-aware AI usage. Our main analyses focus on a within-subject subsample of participants (n = 116) who regularly use both systems, allowing direct comparison of user perceptions and behaviors across the two platforms.
Our primary contributions are:
  1. (1)
    We show that participants indicate greater trust, higher perceived privacy and less experienced LLM hallucinations with their university’s customized LLMaaS chatbot compared to ChatGPT.
  2. (2)
    Drawing on the TrAM framework (Schlicker et al., 2025), we demonstrate that customizations can serve as cues influencing users’ perceptions and behavior toward LLMaaS chatbots.
  3. (3)
    We extend research on calibrated trust by broadening the concept of calibration to include privacy and hallucinations, demonstrating their relevance in an actual academic LLMaaS use-case.
  4. (4)
    We provide organizations and developers with practical recommendations regarding the deployment and customizations of LLMaaS chatbots to support their informed, safe and sustainable use.
2. Theoretical Background and Hypotheses
2.1. LLM (as-a-Service) Chatbots at Universities
LLM chatbots, such as ChatGPT, are widely used in academic settings, with significant adoption among university students (von Garrel and Mayer, 2023; Amani et al., 2023) and researchers (Kapania et al., 2025; Syed et al., 2024; Abdelhafiz et al., 2024). Thus, universities are caught between enabling students and staff to access and use AI in an informed way to leverage the benefits while also protecting data, privacy, rights, and ethical principles. Introducing “their own” LLMs poses a considerable challenge for universities, due to a lack of internal technical abilities and knowledge about the deployment of AI and the high costs of developing and maintaining a sufficient IT infrastructure (Lins et al., 2021).
LLM as-a-Service (LLMaaS) provide a practical solution by offering easy access to pre-trained models via cloud providers (Wang et al., 2024; Syed et al., 2025). This approach is a subset of AI as-a-Service (AIaaS), broadly defined as “cloud-based systems providing on-demand services to organizations and individuals for deploying, developing, training, and managing AI models” (Lins et al., 2021, p. 424). These systems enable organizations to adopt and afford AI technologies without requiring extensive technological expertise (Sundberg and Holmström, 2023; Syed et al., 2025). Moreover, LLMaaS solutions emphasize enhanced data protection and security, offering a viable option for institutions like universities to leverage AI’s advantages while adhering to stringent compliance and security standards (Paavola, 2024).
One proposed advantage of LLMaaS is that organizations can customize various aspects of LLMs to better align with their unique requirements and objectives. In this context, customizations are defined as configurations of parameters at various architectural layers to improve the fit between the system and organization-specific needs (Diaferia et al., 2022). In the broader context of the AI supply chain (Cobbe et al., 2023) such adaptations can occur in a transparent or obscure way for end users (Neumann and Singh, 2025). Table 1 outlines customization options, which encompass customizations of data, models, algorithms, and infrastructure (Diaferia et al., 2022), security and governance adaptations (Chen et al., 2024b), customizations that relate to the design of the user interface and experience, and choices regarding the organizational integration.
Table 1. Examples of customization aspects of AIaaS.
Category Examples References
Model fine-tuning and Configuration Adjust/add system prompts, fine-tuning on proprietary data, hyperparameter configuration, domain-specific knowledge integration, customizable output. (Lins et al., 2021; Diaferia et al., 2022; La Malfa et al., 2023; Hajipour et al., 2023; Parthasarathy et al., 2024; Neumann et al., 2025)
Integration and Interoperability API integration, integration with other tools and systems. (Gan et al., 2023; Lins et al., 2021)
Security and Governance Access control and security, compliance and regulatory requirements, monitoring and evaluation. (Gan et al., 2023; Doosthosseini et al., 2024; Parthasarathy et al., 2024; Iqbal et al., 2024)
User Interface and Experience Incorporation of corporate design, sustainability measures (e.g., token visualization), user profiling and personalization, customizable error handling, multi-language support. (Casati et al., 2019)
Organizational integration Providing context-specific training or workshops, providing context-specific information and use recommendations regarding the model. (Nagarajan et al., 2024)
To date, existing literature has focused on methods and effects of model fine-tuning and configuration, integration and interoperability, as well as security and governance-related consequences of LLMaaS. These aspects are critical because adapting AI systems, for example to ensure data protection or to fine-tune them for internal company data, plays a key role in their secure and human-centered integration into organizations. Accordingly, research and reports on LLMaaS deployment in organizations emphasize primarily technical customizations and their impact on system performance or functionality. For instance, a study on a customized anatomy chatbot showed that tailoring the knowledge base with Retrieval Augmented Generation (RAG) increased factual accuracy (Arun et al., 2024), while reports on the organizational system “FhGenie” highlight aspects such as system architecture and load balancing (Weber et al., 2024). Meanwhile, other customizations pertaining to the user interface of LLMaaS chatbots or their organizational integration – and especially their impact on the end user – receive little attention.
$2.$ Calibration of System Characteristics and User Perceptions: The Role of Cues
Ideally, trust in automated systems should be calibrated, which means that users’ perceived trustworthiness should accurately match the actual capabilities and reliability of the system (Lee and See, 2004). Miscalibration can manifest as either undertrust, leading to system disuse, or overtrust, resulting in system misuse (e. g., failures of monitoring or decision biases) (Parasuraman and Riley, 1997; Wischnewski et al., 2023).
Whether calibrated trust can be achieved depends both on the availability of information about the system and how users perceive and interpret that information (Schlicker et al., 2025). Users’ trust and corresponding use of AI depend on, i. e., “any information element that can be used to make a trust assessment about an agent” (de Visser et al., 2014, p. 253). The TrAM specifies this process, positing that users assess the actual trustworthiness of a system based on available cues, leading to the perceived trustworthiness (Schlicker et al., 2025). The accuracy of this assessment relies on the relevance and availability of cues on the system’s side, as well as the human’s ability to detect and utilize them. On the one hand, actual system characteristics themselves can function as cues, for example information about the training data or the system’s outputs. On the other hand, users may also rely on cues that are less directly tied to a system’s actual qualities, such as interface design or anthropomorphism, which can still strongly influence their perceived trustworthiness (Gong et al., 2024). Such cues may therefore induce unwarranted trust (or distrust) and lead users to perceive systems as more (or less) trustworthy than their actual reliability would justify.
We argue that the concept and relevance of calibration applies not only to trust, but to user perceptions more broadly. For instance, privacy perception calibration can describe the alignment between users’ assumptions about how their data is handled and the system’s actual privacy safeguards. Miscalibrations in this domain have likewise been documented: Windl et al. (2025) showed that misleading cues about browser tracking protection can give users a false sense of privacy. Similarly, hallucination alertness calibration refers to the alignment between users’ cautiousness toward potential incorrect outputs and the system’s actual hallucination tendency. For example, difficulties in detecting biases in LLM outputs have been observed among students in academic use contexts, pointing to a gap between perceived and actual model behavior, a pattern that could similarly occur with hallucinations (Prabhudesai et al., 2025). Accordingly, we propose to conceptualize calibration of user perceptions in the context of LLM use more broadly and understand this as the degree of fit between actual system features (e. g., hallucination tendency, privacy safeguards) and end users’ perceptions of those corresponding system features (e. g., experienced hallucinations or perceived privacy).
Building on the TrAM framework (Schlicker et al., 2025), we focus on the role that customizations of LLMaaS chatbots play as cues that moderate this calibration process between actual and perceived system characteristics. We argue that it is essential to recognize that end users could perceive customizations of LLM chatbots as reflections of the system’s capabilities, goals, and inner workings, even if customizations do not actually affect the functioning of the system. Figure 1 illustrates our proposed interplay between customizations (partially functioning as cues) of the LLMaaS chatbot in relation to actual system characteristics and user perceptions or self-reported behaviors.
Figure 1. System characteristics, customization cues, and user perceptions for the examined university’s customized LLMaaS chatbot. Solid lines represent a positive effect on the associated construct, dotted lines represent a negative effect. Customizations colored in red influence actual system characteristics, while those in blue act as cues expected to affect user perceptions or behavior. Red-and-blue customizations are expected to impact both. Footnotes: ¹Lower temperature may reduce the hallucination tendency by favoring higher-probability tokens and making the output more deterministic (Huang et al., 2025). :
2.3. Trust in AI Systems
2.3.1. Background
Trust, as a key determinant of technology adoption (Van Pinxteren et al., 2019; Choudhury and Shamszare, 2023), has been recognized as an important factor for LLM chatbot adoption among academic staff and students (Bhaskar et al., 2024; Tiwari et al., 2024). Research has shown that both insufficient and excessive trust in AI systems can be critical, as a lack of trust may hinder adoption, while overtrust can lead to risks such as reliance on incorrect or potentially harmful outputs (Buçinca et al., 2021; Wagner et al., 2018). This underscores the importance of ensuring that end users place an appropriate – i. e., calibrated – level of trust in LLM chatbots.
Trust is commonly defined as the trustor’s positive expectations towards the trustee, combined with the trustor’s vulnerability and uncertainty within the specific context (Balayn et al., 2025; Lee and See, 2004). Research has shown that trust levels in human-human interaction depend on three perceived qualities of the trustee: ability, integrity, and benevolence (Mayer et al., 1995). In the context of human-automation interaction, these three factors have been reframed as performance, process (e. g., availability, confidentiality, understandability of a system) and purpose (e. g., the developer’s or deployer’s intentions) (Lee and See, 2004). Additional factors have been identified as particularly important for trust in LLMs. These include transparency (Huang et al., 2023; Schwartz et al., 2023), access to a human operator (Nordheim et al., 2019), perceived privacy and robust data security (Huang et al., 2023; Pawłowska-Nowak, 2024).
2.3.2. Customizations and Hypotheses
Customizations of LLMaaS chatbots that feature the corporate design in the user interface could serve as a salient trust cue for users, increasing the perceived trustworthiness of the system in two dimensions: (a) purpose and (b) process. The perceived purpose of the system may appear more trustworthy when the customizations incorporate university branding (see Figure 2, Label 1) [^1] Figure 2 shows a replica of the LLMaaS chatbot’s user interface, with customizations marked in red., as the university is likely to be seen as a benevolent provider with minimal financial or marketing-driven motives (Nordheim et al., 2019). In this context, university branding refers to the integration of the university’s logo and color scheme into the user interface, creating a visually prominent link between the system and the institution. University branding may also create a sense of familiarity, a factor found to foster interpersonal trust and that might also support trust in AI when users associate it with a familiar institution (Alarcon et al., 2016). Second, the process dimension of the system may be perceived as more trustworthy because it is easily accessible and embedded in the existing organizational framework. Here, features such as the option to contact a human operator for assistance could contribute to users’ trust through perceived openness and accessibility of the system (Lee and See, 2004). Additionally, a customized LLMaaS chatbot may be perceived as more transparent due to context-specific information about the system available on the university’s website. These factors together lead to the hypothesis:
H1: Users report higher levels of trust in the customized LLMaaS chatbot compared to a commercial LLM chatbot.
We further hypothesize that higher trust in the organization’s customized LLMaaS chatbot will likely depend on university members’ trust in the university itself. Organizational trust describes the extent to which individuals trust an organization (Pirson and Malhotra, 2011). Prior research suggests that organizational trust can significantly shape users’ trust in the organization’s system (Balayn et al., 2025; Łapińska et al., 2021). In line with this, Nordheim et al. (2019) have demonstrated that customers’ trust in a company’s service chatbot is influenced by their trust in the company itself. A similar mechanism might apply in the context of academic institutions: users’ trust in a university-provided LLMaaS chatbot may be influenced by their trust in the university. We therefore hypothesize:
H2: The level of trust in the university moderates the effect of system type (customized vs. commercial) on user trust, such that the effect is stronger when organizational trust is higher and weaker when it is lower.
2.4. LLM Hallucinations
2.4.1. Background
When implementing LLMaaS chatbots, universities face the challenge of ensuring that students and staff can use these systems appropriately and securely. A key aspect of this responsibility is addressing the issue of hallucinations. Hallucinations in LLMs describe the generation of content that appears plausible but is factually incorrect or inconsistent with user input (Huang et al., 2025). Hallucinations in human-AI interaction pose a dual challenge: they impair effective collaboration through reduced output quality and undermine users’ trust (Gong et al., 2024; Amaro et al., 2024). In the academic context, hallucinations can pose particular challenges, as they can reduce perceived accuracy and usefulness, increase confusion, and slow task completion (Steinbach et al., 2025). Such issues are compounded by the fact that LLMs have been shown to often fabricate inaccurate scientific references (Aljamaan et al., 2024; Chelli et al., 2024). Despite various efforts and technological advances, hallucinations cannot be completely eliminated from LLMs (Xu et al., 2024; Banerjee et al., 2024). This issue calls for a focus on the human side, with the goal of identifying strategies to help users anticipate, mitigate, and react to hallucinations.
2.4.2. Customizations and Hypotheses
We consider two customization aspects of the user interface: corporate design and warnings about false information. Research has shown that the presence of logos from reputable companies can create positive associations, enhancing the perceived credibility of mobile systems (Lowry et al., 2005). We argue that branding a chatbot with a university logo (see Figure 2, label 1) leads users to interpret the system as university-approved. This could potentially make users less critical and cautious in their interactions with the university-branded system, particularly with regard to the risk of hallucinations. Another notable customization choice of LLMaaS chatbots is the display of warnings about hallucinations or false information. Research suggests that such warnings can help users detect hallucinations more frequently (Nahar et al., 2024). While ChatGPT’s interface includes a disclaimer below the prompt box stating “ChatGPT can make mistakes. Please check important information”, LLMaaS solutions allow such warnings to be adapted, removed, or altered. In our study, the customized LLMaaS chatbot included a hallucination disclaimer only on the login page, with no warning displayed during the actual chat interaction (see Figure 2, label 3). We argue that the absence of a persistent warning may result in users being less mentally primed to consider hallucinations or to critically evaluate the accuracy of the system’s output. The university-branded interface and the lack of a persistent warning may lead users to a) behave less cautiously with respect to potential hallucinations and b) detect fewer hallucinations overall. Based on this, we propose the following hypotheses:
H3: Users report less cautious behavior towards hallucinations when using the customized LLMaaS chatbot compared to a commercial LLM chatbot.
H4: Users report fewer experienced hallucinations when using the customized LLMaaS chatbot compared to a commercial LLM chatbot.
To assess whether actual differences in hallucination tendencies between the systems might contribute to differences in users’ experienced hallucinations, we evaluate both systems using benchmark datasets (see Section 4.4).
2.5. Data Security and Privacy Concerns
2.5.1. Background
End users often express concerns about data security and privacy when interacting with LLMs (Ma et al., 2025b; Ali et al., 2025). Reflecting these concerns, recent research has called for privacy-friendly user interfaces of LLM systems (Li et al., 2024; Song et al., 2025; Windl et al., 2025). Such concerns are grounded in tangible risks, including (unauthorized) data retention (Cobbe and Singh, 2021), where providers may retain and use customer data to enhance their models, inference attacks that aim to extract sensitive information from trained models (Kandpal et al., 2024), and the unintended exposure of confidential data during model training or operation (Chen et al., 2024a). To mitigate these risks, technical customization options in LLMaaS offer promising avenues, playing a pivotal role in enhancing users’ perceptions of privacy and security while fostering trust. These options include user-configurable privacy settings that empower users to control how their data is stored and used (Ashley et al., 2002), transparency mechanisms like privacy explanations (Brunotte et al., 2023), which provide clarity about data handling practices, and the implementation of privacy-preserving techniques such as differential privacy (Abadi et al., 2016) during training and fine-tuning processes.
2.5.2. Customizations and Hypothesis
In the LLMaaS chatbot under study, we identified two system characteristics with potential relevance for users’ privacy perceptions: restricted access to internal networks and the exclusion of chat data from model training (see Figure 1). Beyond these direct technical benefits, we posit that other customization measures – even those without tangible impacts on privacy or security – may foster higher perceived privacy levels. In this context, we again draw on the role of corporate branding in the interface design as a salient customization feature: University branding (see Figure 2, label 1) may enhance users perceptions of data security and privacy, as they may believe that their data will be handled with care and not shared with third parties. Prior research supports this assumption, showing that a strong brand image or high brand credibility can reduce perceived privacy risks and increase users’ willingness to disclose personal information online (Wang, 2019; Myerscough et al., 2008; Jain et al., 2022). Thus, we propose:
H5: Users perceive greater privacy when using the customized LLMaaS chatbot compared to a commercial LLM chatbot.
2.6. Sustainable AI Use
2.6.1. Background
In this final part, we broaden the perspective to considerations that pertain to the global strategy of an organization, specifically the corporate social responsibility (CSR). CSR refers to situations in which organizations pursue strategies that go beyond business needs and legal regulations but aim at furthering social good (McWilliams and Siegel, 2001). This includes, e. g., improving an organization’s sustainability and environmental performance (McWilliams et al., 2006). We argue that customization of LLMaaS chatbots offers organizations opportunities to align AI deployment with their CSR strategy and corresponding sustainability goals, thereby not only shaping the trust in and perception of the systems but also their environmental impact.
The environmental impact of AI systems and LLMs has become an increasingly critical concern in computing research. Recent work by Rafael and Reis (2024) emphasizes how the invisibility of digital energy consumption often leads to unconscious and potentially wasteful usage behavior. This mirrors findings in energy consumption studies (McCarthy et al., 2023) that identify varying levels of user awareness – from “energy unaware” to “energy aware and active” – highlighting how visibility of resource consumption can transform user behavior. Our research extends these insights and considers the impact of customization choices aimed at promoting sustainability-aware AI usage.
2.6.2. Customizations and Hypothesis
Making resource consumption visible has shown promising results across different domains. Penkert and Gröschel (2023) demonstrated that transparency in sustainability information significantly influences user choices, while Sanduleac et al. (2017) found that real-time energy data promotes more responsible consumption patterns. LLMaaS chatbots offer customization options to make AI resource limitations explicit through token visibility (see Figure 2, label 2). The examined LLMaaS chatbot’s interface shows the remaining token contingent for the respective language model as a percentage. We argue that customizing token visibility can foster what we term “sustainability-aware AI use” – a more deliberate approach to AI interaction that considers broader environmental impact. Based on these previous findings, we hypothesize:
H6: Users report higher levels of sustainability-aware AI usage when using the customized LLMaaS chatbots compared to a commercial LLM chatbots.
3. Method
3.1. Study Design
We conducted a quantitative, cross-sectional field study. Data was collected through an online survey during spring 2025 (3rd-12th March) at a major public university that had implemented a customized LLMaaS chatbot at the end of 2024. Participants were recruited via the university chatbot landing page, departmental communications, campus flyers, the university news website, and official university social media channels. Participants who reported regular usage of both chatbots answered the same questions for the university’s customized LLMaaS chatbot and for ChatGPT.
3.2. Description of the University’s Customized LLMaaS Chatbot
The LLMaaS chatbot introduced by the university serves multiple LLMs hosted by different providers. The default model for each chat session, GPT-4o, as well as other OpenAI models, are provided by Microsoft Azure using a datacenter in Sweden. Furthermore, users can choose from a small selection of open-weight models provided by a German AI service center (see Figure 2). The system had 4636 users at the beginning of the data collection period and grew by $20.7\%$ to 5593 users by the end of the period. Customizations regarding the data and model were kept to a minimum, but the university made some adjustments to the user interface and incorporated detailed information and training materials (see Table 2). In contrast to OpenAI’s ChatGPT, the university chatbot supports only text-based interactions and image generation. Other modalities such as audio, document upload, and advanced function calling (e. g., web search) are not supported. Figure 2 shows an anonymized replica of the user interface.
Figure 2. Anonymized replica of the user interface of the university’s customized LLMaaS chatbot. Key elements: (1) university name is displayed; interface follows university’s visual design; (2) remaining monthly token contingent shown in percent; (3) no warning regarding potential hallucinations / incorrect outputs.:
Table 2. Customization features of the studied LLMaaS chatbot.
Category Customization Details
Model fine-tuning and Configuration Temperature1  set to 0 for deterministic outputs. System prompt fixed to: ”You’re a helpful assistant.” for all models, not modifiable by users
Integration and Interoperability Access to multiple models; image generation via DALL-E for OpenAI models
Security and Governance Minimal data collection for service provision; chat data excluded from model training; GDPR-compliant data processing within the EU; access restricted to internal networks (e. g., via VPN)
User Interface and Experience Interface customized with corporate design (e. g., university logo); monthly token contingent visualized as a percentage in the user interface; no warning about hallucinations in bottom section; dropdown menu for choosing active model
Organizational integration Supporting materials provided on the chatbot landing page and intranet; training sessions on effectively using LLMs in higher education
  • 1
    The temperature parameter in LLMs, ranging from 0 to 1, controls response randomness, with values near 0 favoring deterministic outputs and values near 1 producing diverse, creative outputs
$3.$ Ethical Considerations
The study was designed and conducted with great care and attention to ethical criteria. By involving the a) data protection officer, b) work councils for research and administrative staff, c) center for teaching and learning as well as d) information technology department, the rights and interests of all stakeholders were taken into account in the design and implementation of the study. The approval of all parties was sought prior to implementation. The study was also approved by the local ethics committee. Concerning data collection, storage, and sharing, we followed the guidelines of the university’s IT committee and used committee-approved survey software with university-internal data storage, ensuring that only the project team had access to the raw data. The participants were informed about the content and procedure of the study via standard written instructions in line with the guidelines of the German Research Foundation (Schönbrodt et al., 2017). Participation was voluntary, anonymous, and uncompensated.
3.4. Survey Structure and Measures
We designed a comprehensive survey to investigate the perceptions and usage of both the university’s customized LLMaaS chatbot and OpenAI’s ChatGPT. ChatGPT was chosen as the commercial LLM because of its widespread use in Germany (von Garrel and Mayer, 2025). Our data support this choice: in our sample, 333 participants reported using ChatGPT at least once a month, compared to 176 who reported comparable usage of other AI chatbots. During the data collection phase, ChatGPT was powered by OpenAI’s GPT-4o. [^2] See https://help.openai.com/en/articles/6825453-chatgpt-release-notes for a detailed record of ChatGPT’s development during our data collection period. The survey design is summarized in Figure 3. Participants were guided through the survey based on their use/non-use of the chatbots. The main part consists of questions regarding the perception and use of the university chatbot and ChatGPT (green blocks in Figure 3). An overview of all survey questions can be found in the supplementary material (see Section 3.6).
To avoid participant fatigue, all hypothesis-relevant items followed the same response format to minimize cognitive load. We also conducted pilot runs to ensure that average completion time remained under 15 minutes. In addition, only hypothesis-relevant items were mandatory and were placed toward the beginning of the questionnaire to reduce the likelihood of fatigue affecting key measures.
Figure 3. Survey flowchart illustrating participant routing and item blocks based on chatbot usage. Labels at the top of each block indicate the estimated completion time for that section.:
The survey began by capturing usage patterns for both systems: frequency of use (ranging from “never” to “several times a day”) and context of use (6-point scale in response to the item “I tend to use the following AI chatbots for …”, ranging from 1 = “private purposes” to 6 = “study-/work-related purposes”). Next, several aspects of system perception and use were explored: specific use purposes (e. g., information search), perceived usefulness (productivity enhancement, task facilitation), and ease of use (interface clarity, learning curve); the latter two assessed following the Technology Acceptance Model (TAM) (Davis, 1989). For users familiar with either system, we also assessed their awareness of model switching options and use of different large language models available within each platform. For users who have never used or rarely use the university’s LLMaaS, we explored their reasons for non-use or non-regular use.
At the end of the survey, a knowledge test was administered to assess participants’ understanding of the customized LLMaaS chatbot. The test consisted of five items (e. g., “The university chatbot offers models trained by the university” (false)) with response options “true,” “false,” and “I don’t know.” The correct answers of the participants were summed to create a knowledge score ranging from 0 to 5. Participants were debriefed on the following page by being shown the correct answers. At the end of the study, participants could voluntarily report demographic information and were given the opportunity to provide comments in an open text field.
The main part of the survey was structured around the key constructs (see hypotheses). Participants who used both systems responded to the same set of items twice: once for the university’s customized LLMaaS chatbot and once for ChatGPT. Only participants who reported using a system at least once a month were presented with the full set of items about the respective system. Participants were reminded to answer the questionnaire items with regards to their work-related or academic use of the system. We used five-point Likert scales ranging from “strongly disagree” to “strongly agree” for all scales. The questionnaire was available in both German and English, with items from previously published scales translated for the German version. Internal consistencies (Cronbach’s $\alpha$) for the scales are reported in Table B1 and Table B2.
Trust (in the System)
Trust in the two systems was assessed by the facets of global trust, benevolence/purpose, integrity/process, and ability/performance using four items from Wischnewski et al. (2025) (e. g., “I trust [the university chatbot/ChatGPT].”), forming an overall trust measure. We selected this scale for its recent validation, context specificity, and strong fit to human–AI interaction contexts, as well as its theoretical grounding in calibrated trust research (Lee and See, 2004; Parasuraman and Riley, 1997).
Trust in the university
Trust in the university was measured with three items following the trust in management scale by Bøe (2018) (e. g., “My university will safeguard my interests.”). We selected this organizational trust scale due to its strong psychometric validation and its applicability to university contexts. In the validation study, it also predicted intention to continue using institutional technology (Bøe, 2018).
Hallucinations
We investigated users’ cautious behavior toward LLM hallucinations, including their verification practices and task-specific trust decisions, with three items (e. g., “I check important information from [the university chatbot/ChatGPT] responses through external sources.”). Users also reported their experienced hallucinations with three items (e. g., “I have noticed that [the university chatbot/ChatGPT] makes up facts that do not correspond to reality.”). The items were self-developed, as no existing validated scale captured these aspects of LLM hallucinations. Two of the three items measuring experienced hallucinations build on the first two items of the “AI Hallucination Potential” scale by Christensen et al. (2024), which we adapted to our context. We added a third item based on the definition of hallucinations provided by Huang et al. (2025). The items measuring cautious behavior toward hallucinations capture three key user strategies: deciding on LLM use based on the importance of factual accuracy, verifying outputs externally, and adjusting prompts to improve reliability. These strategies are informed by existing research on LLM hallucination mitigation (Barkley and van der Merwe, 2024; Gordon et al., 2023). The internal consistency of the experienced hallucinations scale was acceptable, while cautious behavior toward hallucinations showed low internal consistency. Item-level analysis indicated that removing individual items did not increase reliability beyond $\alpha=0.41$. We discuss this limitation in Section 5.4. Discriminant validity between the two constructs was supported by low correlations, indicating that cautious behavior and experienced hallucinations represent distinct constructs.
Privacy concerns
We measured privacy concerns with three items adapted from Hsu and Lin (2018) (e. g., “My decision to use [the university chatbot/ChatGPT] exposes me to privacy risks.”). To increase content validity and contextual relevance for our study, we added two self-developed items. One of them assesses users’ belief in the confidentiality of the system; the other addresses concerns about chat conversations being read by third parties. This addition was motivated by recent research suggesting that such concerns may be particularly relevant for students using chatbots in academic contexts (Chen et al., 2023).
Sustainability-aware AI usage
We measured the concept of sustainability-aware AI usage with a three-item scale capturing environmentally conscious behavior and energy-related considerations (e. g., “I often think about the energy consumption of [the university chatbot/ChatGPT].”). As no validated instruments for the sustainable use of LLM chatbots currently exist, the items were self-developed. Content validity of the scale was established through systematic consideration of key dimensions of environmentally conscious AI interaction behavior based on literature review (Almanasreh et al., 2019). The items capture three facets of sustainable AI usage: resource-efficient prompting (message length consideration), environmental impact awareness (energy consumption consciousness), and deliberate ecological decision-making (task-specific usage evaluation) (Sanguinetti and Atzori, 2024; Graves et al., 2025).
Since the customized LLMaaS displays the remaining number of tokens as a percentage (see Figure 2, label 2), we additionally assessed the customized LLMaaS’ users’ awareness of token usage with two items (e. g., “During use, I pay attention to how many tokens I still have available.”).
3.5. Participants
A total of $N=526$ participants took part in the survey. A subsample of $n=116$ participants who regularly used both systems was included in the within-subject analyses. We determined our target sample size through an a priori power analysis ($f^{2}=0.05$, $\alpha=.05$, $\beta=.80$), which suggested approximately 250 participants for adequately powered within-subject analyses. The achieved subsample provides sufficient power to detect medium-sized effects, although smaller effects may remain undetected. Within the total sample, 186 participants completed the questionnaire for the university’s chatbot (customized LLMaaS) and 333 for ChatGPT (commercial LLM chatbot). An additional 123 participants reported using neither system at least once per month. Figure A1 shows an overview of participant groups according to their system usage.
Compared to the total sample, the subsample of dual-system users included fewer students and more research and administrative staff (see Table A1 for an overview and comparison of demographic characteristics). This subgroup also had a slightly higher proportion of male participants and was older on average. Within this subsample, $10.3\%$ reported using a paid ChatGPT plan, while $87.1\%$ used the free version (see Table A2).
3.6. Open Science
This study was preregistered on OSF [^3] Non-anonymous preregistration link removed for peer-review. Minor linguistic and structural adjustments were made to the preregistered hypotheses to improve clarity and consistency with the final study design. The dataset and supplementary materials (codebook, questionnaire overview) are available at OSF [^4] https://osf.io/vqsem/files/osfstorage?view_only=53ba0188a6db4a54aa08a1f34bf9078f. We excluded privacy-sensitive entries from the dataset prior to publication.
4. Results
4.1. Descriptive Results: Chatbot Usage Patterns and Knowledge Test
Among the subsample of participants who had used both systems ($n=116$), the university chatbot was most commonly used several times a month ($32.8\%$) or several times a week ($25.0\%$), with only $9.5\%$ using it several times a day. In contrast, ChatGPT had a higher proportion of frequent users, with $19.0\%$ using it several times a day and $28.4\%$ several times a week (see Figure C1 for full distribution). Regarding usage purposes, both systems were most often used for reviewing or improving texts ($70.7\%$). Other prominent use cases included text creation ($54.3\%$ university chatbot; $61.2\%$ ChatGPT) and information search ($52.6\%$ university chatbot; $67.2\%$ ChatGPT; see Figure C2 for a more detailed overview). Awareness of model-switching functionality in the university chatbot was relatively high: $78.4\%$ of dual users knew about the option to change the underlying LLM. However, only $41.4\%$ had actively switched from the default model (GPT-4o) to another. Regarding performance on the knowledge test about the university’s customized LLMaaS, $24.1\%$ answered four questions correctly and $21.6\%$ achieved a perfect score. $14.7\%$ did not answer any item correctly (full distribution see Table C1).
4.2. Hypothesis Testing: Comparisons of the Customized LLMaaS Chatbot vs. ChatGPT
4.2.1. Differences in Perceived Trust, Hallucinations, Privacy and Sustainability-Aware Use
We conducted paired t-tests (Table 3) to test H1 and H3-H6. To account for the increased risk of Type I error due to multiple hypotheses testing, we applied Holm’s step-down correction to the family of six statistical tests (Bender and Stefan Lange, 2001; Holm, 1979). This procedure controls the family-wise error rate (FWER) at $\alpha=0.05$ and was chosen because it is uniformly more powerful than the classical Bonferroni correction while still strictly controlling the FWER (Bender and Stefan Lange, 2001). The six $p$-values were ordered from smallest to largest and sequentially compared against adjusted α-thresholds (i. e., $\alpha$/$k$, where $k$ is the number of remaining tests). The procedure was terminated once a $p$-value exceeded the corresponding adjusted α-threshold; the respective hypothesis and all subsequent ones were rejected. H6, concerning sustainability-aware AI use, was the first hypothesis to exceed its adjusted α-threshold ($p=0.03$ vs. $\alpha=0.0167$), resulting in its rejection along with all subsequent hypotheses in the sequence. Consequently, H3, concerning cautious behavior toward hallucinations ($p=0.19$) and H2 (moderation: system type $\times$ trust in the university; $p=0.44$) were also rejected.
H1, H4 and H5 were supported, as the p-values remained below the respective adjusted $\alpha$-threshold: participants (H1) reported significantly higher levels of trust in the customized LLMaaS compared to ChatGPT, (H4) experienced significantly fewer hallucinations with the customized LLMaaS compared to ChatGPT, and (H5) reported significantly fewer privacy concerns for the customized LLMaaS compared to ChatGPT.
Table 3. Paired t-test comparisons of the customized LLMaaS chatbot (UniGPT) vs. ChatGPT (n = 116), ordered by p-values used for Holm correction.
Dependent Variable M (UniGPT) SD (UniGPT) M (ChatGPT) SD (ChatGPT) t df p Holm $\alpha$1 d2 95% CI3
Privacy concerns (H5) 2.38 0.95 3.61 0.85 -12.42 115 ¡ 0.001 0.0083 -1.15 [-1.39, -0.92]
Trust (H1) 3.60 0.72 3.08 0.79 7.84 115 ¡ 0.001 0.01 0.73 [0.52, 0.93]
Experienced hallucinations (H4) 3.23 1.00 3.85 0.85 -7.76 115 ¡ 0.001 0.0125 -0.72 [-0.92, -0.52]
Sustainability-aware AI use (H6) 2.33 1.18 2.22 1.17 2.16 115 0.03 0.0167 0.20 [0.02, 0.38]
Cautious behavior toward hallucinations (H3) 4.18 0.68 4.13 0.67 1.31 115 0.19 n/a4 0.12 [-0.06, 0.3]
Usefulness5 4.01 0.71 3.98 0.82 0.41 115 0.68 0.04 [-0.14, 0.22]
Ease of use5 4.59 0.65 4.48 0.64 2.12 115 0.04 0.20 [0.01, 0.38]
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below α; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
4.2.2. Moderating Role of Organizational Trust
To test whether trust in the university moderated the system type effect on user trust (H2), we fitted a Linear Mixed-Effects Model (see Table 4) using the lme4 package (Bates et al., 2015). We used a Mixed-Effects Model rather than a simpler linear model (as preregistered) to properly account for the non-independence of the two trust ratings provided by each participant. The model predicted trust from system type (0 = ChatGPT, 1 = customized LLMaaS chatbot), mean-centered trust in the university (to improve interpretability of the intercept), and their interaction, with a random intercept for each participant to capture between-person variability.
Table 4. Mixed-Effects Model of system and trust in the university on user trust (number of observations: 232; $n=116$).
Effect Estimate SE df t p
(Intercept) $3.075\textsuperscript{***}$ $0.066$ $184.7$ $46.33$ <0.001<0.001
System $528\textsuperscript{***} 0.0$670.067 $114.0$ $7.83$ <0.001<0.001
TrustInUniversity_centered $272\textsuperscript{**} 0.0$840.084 $184.7$ $3.25$ $0.001$
System:TrustInUniversity_centered $0.066$ $0.085$ $114.0$ $0.78$ $0.440$
Model fit
AIC 493.72
BIC 514.40
Log Likelihood -240.86
Random effects
Num. groups: Bogen 116
Var: Bogen (Intercept) 0.25
Var: Residual 0.26
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$
H2 was not supported, as the interaction between system type and trust in the university was not significant. This indicates that the positive relationship between trust in the university and system trust did not differ between ChatGPT and the university’s customized LLMaaS.
4.3. Controlling for Confounding Variables
We extended our preregistered analyses to account for potential confounding variables. In this section, we report robustness checks that address factors which could plausibly influence our main hypothesis tests, while the following section covers exploratory analyses aimed at uncovering additional patterns beyond our preregistered scope.
4.3.1. Usage Frequency does not moderate the relationship between system type and experienced hallucinations
To test whether the difference in experienced hallucinations between the customized LLMaaS chatbot and ChatGPT was moderated by usage frequency, we estimated a Cumulative Logit Model (Table 5) including both main effects and their interaction using the VGAM package in R (Yee, 2010). Usage frequency did not significantly moderate the relationship between system type and experienced hallucinations, as indicated by a Wald test of all interaction terms ($\chi^{2}=9.60$, $df=5$, $p=0.09$). We chose this modeling approach because the usage frequency was measured with an ordinal scale. We accordingly transformed hallucination ratings into five ordered categories using equally spaced cut points. Among the usage frequency contrasts, the lowest p-value was observed for the fifth-order polynomial, which tests whether hallucination reports follow a complex pattern across usage levels. However, this effect did not fall below the alpha level of $0.05$ and was therefore not statistically significant.
Table 5. Cumulative Logit Model: System type and usage frequency on experienced hallucinations (number of observations: 227).
Effect Estimate SE z p
System (ChatGPT vs. UniGPT) $0.955\textsuperscript{**}$ $0.309$ 3.09-3.09 $0.002$
Frequency (Linear) 0.613-0.613 $0.539$ 1.14-1.14 $0.255$
Frequency (Quadratic) $0.004$ $0.472$ $0.01$ $0.994$
Frequency (Cubic) 0.511-0.511 $0.556$ 0.92-0.92 $0.358$
Frequency (4th order) 0.095-0.095 $0.555$ 0.17-0.17 $0.864$
Frequency (5th order) 0.811-0.811 $0.452$ 1.79-1.79 $0.073$
System × Frequency (Linear) $0.954$ $0.743$ $1.28$ $0.199$
System × Frequency (Quadratic) 0.912-0.912 $0.653$ 1.40-1.40 $0.163$
System × Frequency (Cubic) 0.184-0.184 $0.809$ 0.23-0.23 $0.821$
System × Frequency (4th order) 1.158-1.158 $0.832$ 1.39-1.39 $0.164$
System × Frequency (5th order) $0.171$ $0.683$ $0.25$ $0.802$
Model fit
Residual Deviance 608.20
Log Likelihood -304.10
AIC 638.20
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$
Model estimated on 227 observations (df = 893)
4.3.2. Model switching does not affect system differences
To examine whether switching the default language model (GPT-4o) within the university’s customized chatbot influenced the results, we repeated the within-subject comparisons using paired-samples t-tests on a reduced subsample of participants who reported never having changed the underlying model ($n=68$; see Appendix D for detailed test statistics). The pattern of results remained identical: the same differences between systems reached statistical significance (or not), and all observed effect directions were consistent with those in the full sample. Consequently, the interpretation of our findings and conclusions regarding the tested hypotheses remained unaffected. We therefore report the results based on the full sample of users who had experience with both systems ($n=116$) in the main text.
4.3.3. Prior knowledge does not affect system differences
To control for potential confounding effects of participants’ knowledge about the unversity’s customized LLMaaS chatbot, we included knowledge scores as a covariate in our analysis. The system differences proved robust to the prior knowledge of the participants, as neither the knowledge scores nor their interaction with the system type significantly predicted any outcome measures (all $p>0.11$; see Table D1 for detailed results). To examine this, we fitted Linear Mixed-Effects Models using the lme4 package, specifying fixed effects for system type, knowledge scores, and their interaction, as well as random intercepts for participants.
$4.$ Exploratory Analyses
4.4.1. Benchmark evaluations reveal more hallucination generation but better hallucination recognition for the customized LLMaaS chatbot
To complement the self-reported experienced hallucinations of users, we conducted two exploratory benchmark evaluations to assess the hallucination tendencies of the chatbots: TruthfulQA and HaluEva. The university’s customized LLMaaS chatbot and ChatGPT were tested using the same LLM, namely GPT-4o.
TruthfulQA
TruthfulQA (Lin et al., 2022) is a benchmark designed to evaluate whether language models generate factually accurate responses to questions that are prone to false answers due to common misconceptions or false beliefs. A model performs well on this benchmark if it avoids reproducing widespread falsehoods learned from human-generated texts. We evaluated both systems on 50 questions from the TruthfulQA dataset. We selected 19 questions from the “Education & Science” category to reflect the university context of our study, and sampled 31 additional questions from the remaining categories. The university’s customized LLMaaS chatbot produced hallucinated answers in 25 out of 50 cases ($50\%$), whereas ChatGPT hallucinated in 19 out of 50 cases ($38\%$). Notably, the university chatbot hallucinated more frequently across both the domain-specific and general subsets of the benchmark.
HaluEval
HaluEval (Li et al., 2023) is a large-scale collection of generated and human-annotated hallucinated samples for evaluating the performance of LLMs in recognizing hallucinations. We tested both systems on $10,000$ QA examples from the HaluEval dataset. Results showed that both systems achieved comparable performance, with the university’s LLMaaS chatbot reaching slightly higher values than ChatGPT across all metrics. Specifically, it achieved higher accuracy ($0.7190$ vs. $0.7011$), precision ($0.6671$ vs. $0.6564$), recall ($0.8807$ vs. $0.8472$), and F1-score ($0.7591$ vs. $0.7397$).
4.4.2. Perceived usefulness does not differ, ease of use is greater for the customized LLMaaS chatbot
For participants who had used both systems, paired t-tests were conducted to compare perceived usefulness and ease of use (see Table 3). For usefulness, no statistically significant difference was found between the customized LLMaaS ($M=4.01$, $SD=0.71$) and ChatGPT ($M=3.98$, $SD=0.82$), $t(115)=0.41$, $p=0.68$, $d=0.04$. For ease of use, ratings were statistically significantly higher for the customized LLMaaS ($M=4.59$, $SD=0.65$) compared to ChatGPT ($M=4.48$, $SD=0.64$), $t(115)=2.12$, $p=0.04$, $d=0.20$.
4.4.3. Between-subjects comparisons mostly align with within-subjects results
To complement the within-subject analyses, we conducted exploratory between-subject comparisons using Welch’s two-sample t-tests. These analyses included participants who reported regular use of only one system – either the university’s customized LLMaaS chatbot ($n=70$) or ChatGPT ($n=217$). The full set of results is provided in Table E1; overall, the between-subjects results mostly aligned with the within-subjects results, and here we report only findings that deviated. Notably, participants who exclusively used the customized LLMaaS chatbot reported significantly higher sustainability-aware AI usage compared to those who only used ChatGPT. In contrast, the significant within-subject difference in trust in favor of the customized LLMaaS chatbot was not replicated in the between-subjects comparison. Similarly, the significant within-subject effect for perceived ease of use (favoring the customized LLMaaS chatbot) was not statistically significant in the between-subjects analysis. All other outcomes, including privacy concerns, experienced hallucinations, cautious behavior toward hallucinations and perceived usefulness, showed consistent patterns across both analysis approaches. Given that ChatGPT Pro users are not subject to token limits, which may compromise the comparability of groups for sustainability-aware AI usage, we conducted an additional comparison between exclusive ChatGPT Free users ($n=176$) and users of the customized LLMaaS chatbot ($n=70$). Detailed results are provided in Table E2. This analysis showed the same pattern of results as the full-sample between-subjects comparison, including significantly higher sustainability-aware AI usage among customized LLMaaS users and no significant difference in trust.
4.4.4. Customized LLMaaS chatbot is used mainly for academic purposes, ChatGPT for both academic and private purposes
Participants were instructed to respond to all dependent variable items with regard to their study- or work-related use of the AI systems in the context of their everyday university life. However,we also report on the overall usage tendencies that participants indicated across the private vs. work/study usage context. Participants who used both systems reported using the customized LLMaaS chatbot with a strong orientation toward study- or work-related contexts ($M=5.48$, $SD=0.81$; scale range: 161-6), whereas reported usage of ChatGPT was more balanced across the spectrum ($M=3.32$, $SD=1.61$). A paired samples t-test confirmed that this difference was statistically significant, $t(115)=13.36$, $p<.001$, $d=1.24$. Notably, $64.7\%$ of participants selected the highest possible work/study-related option (6) for the customized LLMaaS chatbot, while only $11.2\%$ did so for ChatGPT.
5. Discussion
This study examined how customizations of LLMaaS chatbots affect users’ trust, perception and usage of such tools in the context of everyday university life. Our findings support the hypothesized differences in users’ perceptions of the customized LLMaaS chatbot versus ChatGPT in regard to trust, privacy and experienced hallucinations. In the following, we interpret the results and elaborate on challenges in achieving calibrated user perceptions of LLMaaS. Finally, we discuss the theoretical and practical implications, as well as the limitations, of our study.
5.1. Understanding User Perceptions of a Customized LLMaaS Chatbot vs. ChatGPT
Participants showed higher trust in the university’s customized LLMaaS chatbot compared to ChatGPT (H1). This finding supports the assumption that users draw on cues about the system such as the corporate interface design in their trust assessments. It also aligns with the underlying theory of trust in automation (Lee and See, 2004), which highlights process and purpose as key dimensions of user trust. Our findings support the relevance of these dimensions in shaping trust in LLM chatbots, as the customized LLMaaS chatbot may have been perceived as more easily accessible and integrated into institutional structures (process), and as provided by a benevolent, responsible actor (purpose), thereby fostering greater trust. The result further supports our rationale that customizations may enhance user trust by increasing factors such as perceived openness or transparency of the system, which is consistent with previous findings (Schwartz et al., 2023; Nordheim et al., 2019). Moreover, it suggests that the role of familiarity in fostering interpersonal trust (Alarcon et al., 2016) might also be relevant in human–chatbot interactions.
The hypothesized moderation effect of trust in the university (H2) was not supported, i. e., the effect of system type on user trust did not depend on how much users trusted the university. Instead, both the system type (customized LLMaaS chatbot) and higher trust in the university independently contributed to higher trust in the system. This finding may reflect the influence of users’ general propensity to trust, that is, a dispositional tendency to attribute trustworthiness across different social and technological entities (Mcknight et al., 2011; Heyns and Rothmann, 2015). Overall, the results point to additive rather than interactive effects of system type and trust in the university on users trust in the system.
We found no statistically significant difference in how cautious participants behaved toward hallucinations between the customized LLMaaS chatbot and ChatGPT (H3). However, participants reported more experienced hallucinations with ChatGPT than with the customized LLMaaS chatbot (H4). This combination of results is noteworthy as we expected that the same customizations of the LLMaaS chatbot, the absence of an in-chat hallucination warning and the corporate design, would jointly reduce users’ alertness toward hallucinations, thereby leading to both less cautious behavior and fewer reported hallucinations. However, our findings only support the latter. While the lack of an effect for H3 could be due to the low internal consistency of our self-developed scale for cautious behavior, we discuss further possible interpretations. One possible explanation for this pattern is that interface customizations may have reduced users’ mental alertness only to a limited degree. This may have influenced the detection of hallucinations, which rely more on automatic or intuitive processes. However, it may not have affected more reflective, effortful strategies for preventing hallucinations or verifying outputs. These interpretations gain further nuance when considered alongside our exploratory benchmark evaluations. TruthfulQA directly assessed hallucination generation and indicated that the LLMaaS chatbot in fact produced more hallucinated answers than ChatGPT. In contrast, HaluEval evaluated hallucination detection performance and showed that the university chatbot slightly outperformed ChatGPT. Taken together, these results suggest a miscalibration between users’ perceptions and objective performance: although participants perceived the customized chatbot as less hallucination-prone, benchmark evidence indicates that it may generate hallucinations more frequently, while at the same time showing somewhat stronger abilities in recognizing them. This discrepancy underscores the role of trust-inducing cues and institutional framing in shaping users’ vigilance toward hallucinations, potentially leading to an underestimation of risks.
Higher privacy perceptions were observed for the customized LLMaaS chatbot compared to ChatGPT (H5). These results relate to actual system differences in terms of privacy safeguards (see Figure 1). Additionally, users might have associated the university-hosted system with more responsible or secure data practices due to its institutional branding in the interface design or login-based access within the university network. Prior work has shown that strong corporate branding can reduce perceived privacy risks and increase users’ willingness to disclose personal information online (Wang, 2019; Myerscough et al., 2008; Jain et al., 2022). In line with this reasoning, the university’s institutional branding may have served as a trust signal, e. g., leading users to perceive the system as more accountable and less likely to monetize their data. These findings highlight the role of contextual design cues in shaping perceived privacy and suggest that institutional trust signals may function as privacy heuristics in AI-mediated interaction.
Participants did not report more sustainability-aware AI usage when interacting with the customized LLMaaS chatbot compared to ChatGPT (H6). This finding stands in contrast to previous work showing that making resource consumption visible can foster more sustainable behavior (Penkert and Gröschel, 2023; Sanduleac et al., 2017). Notably, the exploratory between-subjects comparisons showed different results: users who exclusively used the customized LLMaaS chatbot reported more sustainability-aware AI usage than those who exclusively used ChatGPT (this finding also held true when comparing LLMaaS-exclusive users to users of the free, token-restricted ChatGPT version). Interestingly, LLMaaS-exclusive users also reported higher sustainability-aware AI usage than the group of users of both systems showed for either the LLMaaS chatbot or ChatGPT. This pattern could reflect underlying differences between user groups, such that individuals who use only the customized LLMaaS chatbot may differ systematically in values (Shang et al., 2023) or usage motivations (Wunderlich et al., 2019) that relate to more sustainability-aware AI behavior. Alternatively, lower sustainability-aware AI usage associated with ChatGPT might originate from characteristics of the system itself (e. g., no token usage visualization; low-barrier access across devices, including as a mobile app) and carry over to other systems used by the same person. While these interpretations remain speculative, they highlight a promising avenue for future research on how individual traits or system characteristics may shape sustainability-aware AI usage.
Overall, our results support the assumption that customization-related process and purpose cues enhance user trust, and that institutional framing contributes to privacy perceptions. However, they also raise questions about whether such cues affect deliberate behaviors, such as cautious behavior toward hallucinations. The lack of moderation by institutional trust and the mixed results on sustainability-aware AI use suggest that, alongside customization effects, individual traits – such as the propensity to trust or sustainability-related values – might represent a complementary factor that is worth examining in future work on human–AI interaction.
5.2. How Calibrated Are Participants’ Perceptions of the Chatbots?
The concept of calibration has previously been considered primarily in relation to trust (Parasuraman and Riley, 1997; Lee and See, 2004). In this study, we conceptualize the concept of calibration more broadly and build on the TrAM model (Schlicker et al., 2025) by linking actual system features with perceived system characteristics via cues experienced by the user. In general, we see that users displayed higher trust, perceived greater privacy and experienced less hallucinations with the customized LLMaaS chatbot. At first glance, this suggests that users are able to calibrate their perceptions towards an LLM chatbot, as, e. g., the university’s customized chatbot has stronger privacy safeguards and users have less concerns. However, we question whether the extent of the perceived difference is justified and if users’ consequent behavior is “right”. First, we question how well users understand the differences between the systems, given the results of the knowledge test (see Section 4.3). Second, an appropriate calibrated perception is further questioned by the results of the hallucination benchmark evaluations, which allowed us to relate users’ subjective experiences to objective performance metrics under standardized conditions. The results showed that, although participants perceived fewer hallucinations with the customized LLMaaS, objective metrics suggest that this system was at least equally, if not more, prone to hallucinations. Accordingly, the anticipated link between the system’s lower temperature setting and reduced hallucination tendency (see Figure 1) did not hold. In sum, this divergence highlights a critical gap in users’ hallucination alertness calibration, that is, their ability to align vigilance against hallucinations with the system’s actual risk. Users may underestimate the risk of hallucinations due to trust-inducing cues such as branding, institutional context, and lack of persistent warnings, potentially compromising critical engagement with LLM outputs. When this finding is considered alongside the significantly higher perceived privacy in LLMaaS, the question arises as to whether users are overtrusting the customized LLMaaS chatbot.
The results of this study advance the concept of calibrated trust by extending it to broader calibrated system perceptions, including privacy perception calibration and hallucination alertness calibration. By integrating the TrAM framework, we have demonstrated the importance of providing users with accurate system information, particularly in relation to customizations, via cues to enable calibrated system perceptions that support safe and appropriate use.
5.3. Implications
5.3.1. Theoretical Implications
Calibrated user perceptions and the role of cues
Our study adds to the mixed findings on calibrated trust (Wischnewski et al., 2023), by showing that customizations of an LLMaaS chatbot influence users’ trust significantly, even when their knowledge of actual system differences is limited. Furthermore, we extended the concept of calibration to LLM hallucinations and privacy, highlighting the significance of calibrated AI perceptions and use in other aspects. We also extended the TrAM framework (Schlicker et al., 2025) by applying it to users’ perceived privacy and self-reported experienced hallucinations. Our results indicate that customizations of LLMaaS chatbots are perceived by users in the form of cues and can have a significant impact on their perceived privacy, trust in the system and experienced hallucinations. These findings suggest that the TrAM is particularly useful for identifying points in the user experience where misalignments between users’ perceptions and the system’s actual characteristics may occur. This makes it a valuable framework for scholars and practitioners to use when identifying or refining existing cues, or when introducing new ones, in order to foster more calibrated user perceptions.
Our findings also open up important avenues for future research on the cues underlying user judgments of AI systems. While we showed that system customizations influence user perceptions of trust, privacy, and hallucinations, it remains an open question which specific cues users rely on most when forming such judgments. A key next step for theory development in the area of calibrated AI use is to investigate what kinds of cues users notice, which ones they consider most relevant, and how well these cues reflect the actual characteristics of the system. Future studies should seek to systematically compare different cues in terms of their influence on user perceptions and their alignment with actual system properties in order to better understand how (mis)calibrations emerge in practice.
New Directions for User Hallucination Detection Research
This work stimulates new directions for research on the users’ detection of LLM hallucinations, particularly regarding usage context and task characteristics. Prior work has mainly focused on individual-level factors (Dang and Nguyen, 2025), explanation formats (Reinhard et al., 2025), and information characteristics (Tao et al., 2024), but has paid less attention to how system use itself shapes hallucination detection. Our results point to notable differences in the broader usage context of the two systems. Although participants were instructed to answer the questionnaire items with regard to their work- or study-related use, the reported overall usage patterns diverged substantially: within the subsample of both systems users, the customized LLMaaS chatbot was primarily used for academic or professional tasks, whereas ChatGPT use was more evenly distributed between private and academic contexts (see Section 4.4.4). It remains an open question whether such contextual differences affect users’ expectations, vigilance, or standards for verifying outputs, and consequently their likelihood of noticing hallucinations. In addition to context, future research should also consider how different usage purposes influence hallucination alertness and detection. Beyond broad categories such as information search or text creation, more nuanced dimensions of purpose, such as task complexity, topical domain, or the type of user goal (e. g., idea generation vs. factual information retrieval), might influence how carefully outputs are evaluated. Addressing these aspects through controlled experiments could enrich existing research and provide a more comprehensive understanding of hallucination detection.
5.3.2. Practical Implications
Our study provides valuable insights for organizations considering the adaption of a customized LLMaaS chatbot and those actively running one.
Supporting calibrated user perceptions
Most importantly, we advise institutions deploying LLMaaS chatbots to pay attention to the potential discrepancy between the actual and perceived trustworthiness of the system, i. e., ensuring that user perceptions are calibrated correctly in relation to the characteristics of the system. In particular, developers and deployers should ensure that surface-level cues do not create a false impression of the system’s capabilities or privacy safeguards. This could involve actively communicating residual risks, providing context-specific warnings, and offering transparency tools to foster better calibration. The results of our study showed that although a system might not hallucinate less, it can be perceived to do so. Therefore, we strongly suggest implementing a hallucination warning in the user interface. Previous research has shown that warnings can increase hallucination detection without significantly affecting perceived truthfulness of genuine content (Nahar et al., 2024). Similarly, Bo et al. (2025) found that adding a persistent disclaimer below the LLM output was the most effective at improving users’ appropriate reliance in a complex problem-solving task (“Remember to verify this information and adapt it to your specific problem as I may not have complete information about your context.”). Building on this, additional warnings could be provided when the system offers links, encouraging users to verify their accuracy and relevance to their request before using them for further purposes, e. g., in academic writing. Moreover, confidence scores presented alongside generated output and links connecting system responses to the source documents used to generate them could be effective transparency tools to support calibrated user perceptions (Leiser et al., 2023; Reinhard et al., 2025).
At an institutional level, measures could include a mandatory online course providing information on the risks, limitations, and particularities of the system, which users would have to complete to activate their account for using LLMaaS chatbots. One helpful customization that could support the calibrated use of a university’s LLMaaS chatbot would be to integrate a dedicated “audit-assist” mode with structured, example-driven scaffolding features that guide users through systematic audits of model behavior. Such scaffolding, as shown in a recent study (Prabhudesai et al., 2025), can improve students’ critical engagement with LLMs, including their ability to identify biases. Furthermore, alerts could be added to the interface when users ask for medical, legal or financial advice, or when they enter personal data (e. g., names, email addresses, etc.), to facilitate critical output evaluation in high-risk contexts. To encourage appropriate user behavior particularly in relation to privacy, one option for developers of LLMaaS chatbots would be to integrate user-led data minimization tools that help users detect and remove personal information from their prompts (Zhou et al., 2025).
Another promising institutional measure is the integration of usage analytics and feedback loops. By collecting anonymized data on chatbot interactions (e. g., hallucination reports, privacy concerns, trust difficulties) and combining them with periodic user feedback, universities can continuously monitor risks and adapt training or interface features accordingly. Such human-in-the-loop monitoring can help calibrate user perceptions with actual system performance, ensuring that trust remains both appropriate and evidence-based.
Supporting sustainable AI usage
Moving on, organizations could also use customizations to support sustainable AI usage. Although our within-subject analysis did not support H6 regarding sustainability-aware AI use, the significant between-subject difference suggests that customization features, such as displaying the token contingent in the interface, may encourage more sustainable usage behavior. Institutions might leverage such design choices to align AI usage with broader sustainability goals. The token contingent displayed in the chatbot’s user interface is one promising option, while further customizations could include intuitive labels indicating resource efficiency or environmental impact of different models, such as care-label-inspired badges with color-coded usage indicators for energy consumption and runtime (Fischer et al., 2024, 2025).
User acceptance as a baseline
For decision makers in organizations, our findings indicate that introducing customized LLMaaS chatbots can build on a solid baseline of user acceptance. Some customizations may even improve perceived ease of use compared to established commercial LLM chatbots. The exploratory analyses showed that, while users found both systems equally useful, they indicated higher ease of use for the customized LLMaaS chatbot compared to ChatGPT. These findings are valuable in terms of the applicability of customized LLMaaS chatbots in real-world institutional contexts, given that these two variables are key factors in predicting user acceptance of technologies (Davis, 1989), a link that has also been shown for AI systems (Choung et al., 2022; Kelly et al., 2023).
5.4. Limitations
The study has limitations that need to be taken into account in the interpretation of the results and suggest avenues for future research.
First, our main analyses were based on a within-subjects subsample of users who had experience with both systems. While this group enables controlled comparisons, it may also introduce a self-selection bias. Although key effects (e. g., privacy and hallucinations) were replicated in the between-subjects analysis, sustainability-aware AI usage was only significant in the between-subjects comparison. Future studies should therefore examine the generalizability of our findings across user groups (e. g., students vs. employees; single-system vs. multi-system users).
Second, we did not control for potential order effects in the questionnaire. Due to the limited configuration options on the university’s internal survey platform, the university’s LLMaaS chatbot was always evaluated before ChatGPT. While this ordering was consistent for all participants and unlikely to explain system-specific effects alone, it remains a potential confound. Future studies should employ counterbalanced survey designs to rule out such effects.
Third, our study relied exclusively on self-report measures. Future studies should incorporate behavioral data, such as prompt input, output verification actions, or system usage logs, to complement self-reports. Furthermore, the internal consistency of our newly developed scale measuring cautious behaviour towards hallucinations was low. This may be due to heterogeneity in the construct, as it includes both proactive (e. g., adjusting prompts) and reactive (e. g., checking outputs) strategies. Future research should consider splitting and extending this scale to better distinguish between these dimensions and improve psychometric quality.
Fourth, although we conducted a priori power analyses, our within-subjects sample (n = 116) fell short of the planned size (n = 250). While sufficient for detecting medium effects, smaller effects, such as differences in cautious behavior or sustainability-aware AI use, may have gone undetected.
Fifth, the study design does not allow for causal interpretations. While one key strength of our study is that we surveyed real users in a naturalistic setting, thereby generating findings with high ecological validity, this comes at a cost of lower internal validity. This could be addressed in future research through the use of controlled experimental designs to isolate the effects of individual features, while longitudinal studies may shed light on how user perceptions evolve over time. Such designs could also examine potential causal links between specific user perceptions, e. g., user trust and experienced hallucinations.
Sixth, the systems differed not only in branding and context, but also in available features, e. g., ChatGPT supports voice input, web search, and file uploads, which the university chatbot does not. While we did not focus on these differences, they may have influenced user perceptions such as ease of use or perceived privacy. Future research should examine how specific features shape system evaluations.
Finally, we see potential for qualitative follow-up research to better understand how and why the patterns observed in our quantitative findings emerge. Several open-ended comments from participants hint at meaningful underlying reasoning and motivations, e. g.: “I always use [the university chatbot] when I have concerns about data privacy regarding the content submitted to ChatGPT.” Future studies could conduct in-depth interviews with users to explore the use of cues in system assessments and decision-making strategies when choosing between different LLM chatbots.
$5.$ Conclusion
We examined how user-salient customizations in LLMaaS chatbots, such as institutional branding and interface changes, shape user perceptions and behavior. Participants reported greater trust, fewer perceived hallucinations, and lower privacy concerns when using the university’s customized chatbot compared to ChatGPT – despite both systems being based on the same underlying LLM technology. These differences highlight the influence of contextual and interface-level cues on users’ system assessments. Our findings emphasize the importance of calibrated user perceptions not only with regard to trustworthiness, but also in relation to privacy and hallucination risk. In doing so, our work contributes to the TrAM and the concept of calibrated trust, expanding both to include more facets of human-AI interaction. Practically, we provide evidence-based guidance on AI implementation. The careful design of interfaces, features, and integration strategies for LLMaaS chatbots can promote their appropriate use in line with their risks and limitations.
We encourage future research to explore which specific cues users rely on when forming their perceptions and how such alignments affect downstream behavior. Our work therefore opens new research directions to investigate how customization aspects impact users.
Acknowledgements.
We would like to thank Markus Pauly, Nico Föge, and Daniel Klippert for their valuable support with the statistical analyses. We are also grateful to Lin Kyi, Gabriel Lima, and Yixin Zou for their helpful feedback. Our sincere thanks go to all participants for contributing their time and insights, as well as to the involved university stakeholders for their continuous support in enabling this study. This work was supported by the Sponsor Research Center Trustworthy Data Science and Security https://rc-trust.ai (https://rc-trust.ai), one of the Research Alliance Centers within the UA Ruhr (https://uaruhr.de).
Appendix
Appendix A Demographics
Figure A1. Percentage of participants using different chatbot systems.:
Table A1. Demographics comparison: total sample vs. subsample (both systems users)
Total Sample (N = 526) Subsample (n = 116)
$n$ (%)
User Group
Students 262 (49.9%) 22 (19.1%)
Research staff 146 (27.8%) 52 (45.2%)
Administrative staff 89 (17.0%) 35 (30.4%)
Professors 21 (4.0%) 4 (3.5%)
Other 7 (1.3%) 2 (1.7%)
Gender
Male 250 (47.5%) 63 (54.3%)
Female 246 (46.8%) 46 (39.7%)
Diverse 10 (1.9%) 1 (0.9%)
Other 1 (0.2%) 0 (0.0%)
Prefer not to say 18 (3.4%) 6 (5.2%)
Age
Under 25 152 (28.9%) 14 (12.1%)
25–34 219 (41.6%) 53 (45.7%)
35–44 80 (15.2%) 31 (26.7%)
45–54 34 (6.5%) 9 (7.8%)
55–64 38 (7.2%) 9 (7.8%)
65 or older 2 (0.4%) 0 (0.0%)
Table A2. ChatGPT plan usage: All ChatGPT users vs. subsample (both systems users).
All ChatGPT Users (n = 333) Subsample (n = 116)
$n$ (%)
ChatGPT Plan
Paid version 53 (15.9%) 12 (10.3%)
Free version 277 (83.2%) 101 (87.1%)
Not specified 3 (0.9%) 3 (2.6%)
Appendix B Correlation Tables
Table B1. Means, standard deviations, Cronbach’s $\alpha$, and correlations with confidence intervals for the customized LLMaaS chatbot (subsample of both systems users, $n=116$).
Variable M SD $\alpha$ 1 2 3 4 5 6 7
1. Trust 3.60 0.72 0.59
2. Cautious behavior toward hallucinations 4.18 0.68 0.40 -0.06 [-0.24,0.13]
3. Experienced hallucinations 3.23 1.00 0.82 -0.36** [-0.51,-0.19] 0.18 [-0.01,0.35]
4. Privacy concerns 2.38 0.95 0.89 -0.48** [-0.61,-0.32] -0.05 [-0.23,0.13] 0.31** [0.13,0.47]
5. Sustainability-aware AI use 2.33 1.18 0.79 0.03 [-0.16,0.21] 0.04 [-0.14,0.22] -0.01 [-0.19,0.17] 0.21* [0.03,0.38]
6. Perceived usefulness 4.01 0.71 0.86 0.46** [0.31,0.59] -0.07 [-0.25,0.11] -0.19* [-0.36,0.00] -0.28** [-0.44,-0.10] -0.03 [-0.21,0.15]
7. Perceived ease of use 4.59 0.65 0.91 0.33** [0.15,0.48] -0.00 [-0.19,0.18] -0.03 [-0.21,0.15] -0.23* [-0.40,-0.05] -0.05 [-0.23,0.14] 0.28** [0.10,0.44]
8. Trust in the university 3.91 0.80 0.90 0.38** [0.21,0.52] -0.04 [-0.22,0.14] -0.21* [-0.37,-0.02] -0.19* [-0.36,-0.01] -0.00 [-0.19,0.18] 0.33** [0.15,0.48] 0.23* [0.05,0.40]
Note. $\alpha$ = Cronbach’s Alpha. Values in brackets = 95% CI. * $p<.05$, ** $p<.01$.
Table B2. Means, standard deviations, Cronbach’s $\alpha$, and correlations with confidence intervals for ChatGPT (subsample of both systems users, $n=116$).
Variable M SD $\alpha$ 1 2 3 4 5 6 7
1. Trust 3.08 0.79 0.71
2. Cautious behavior toward hallucinations 4.13 0.67 0.40 -0.05 [-0.23,0.13]
3. Experienced hallucinations 3.85 0.85 0.79 -0.23* [-0.39,-0.05] 0.32** [0.15,0.48]
4. Privacy concerns 3.61 0.85 0.84 -0.35** [-0.50,-0.18] 0.00 [-0.18,0.18] 0.29** [0.12,0.45]
5. Sustainability-aware AI use 2.22 1.17 0.75 0.10 [-0.08,0.28] -0.02 [-0.20,0.17] -0.08 [-0.26,0.10] 0.02 [-0.16,0.20]
6. Perceived usefulness 3.98 0.82 0.87 0.56** [0.42,0.67] 0.07 [-0.11,0.25] 0.02 [-0.17,0.20] -0.19* [-0.36,-0.01] 0.00 [-0.18,0.19]
7. Perceived ease of use 4.48 0.64 0.90 0.18 [-0.00,0.35] 0.17 [-0.01,0.34] 0.26** [0.08,0.42] 0.04 [-0.14,0.22] -0.04 [-0.23,0.14] 0.39** [0.22,0.53]
8. Trust in the university 3.91 0.80 0.90 0.27** [0.10,0.44] 0.03 [-0.15,0.21] -0.11 [-0.29,0.07] -0.00 [-0.18,0.18] 0.03 [-0.16,0.21] 0.15 [-0.04,0.32] 0.15 [-0.03,0.33]
Note. $\alpha$ = Cronbach’s Alpha. Values in brackets = 95% CI. * $p<.05$, ** $p<.01$.
Appendix C Descriptive Results
Figure C1. Usage frequency of the university’s customized LLMaaS chatbot vs. ChatGPT in the subsample of both systems users ($n=116$).:
Figure C2. Purpose of use for the university’s customized LLMaaS chatbot vs. ChatGPT in the subsample of both systems users ($n=116$).:
Table C1. Distribution of knowledge test scores for the subsample of both systems users ($n=116$). Scores represent the number of questions answered correctly.
Knowledge Score (0–5) $n$ (%)
0 17 (14.7%)
1 11 (9.5%)
2 22 (19.0%)
3 13 (11.2%)
4 28 (24.1%)
5 25 (21.6%)
Appendix D Controlling for Confounding Variables
[htb] Paired t-test comparisons of the customized LLMaaS chatbot (UniGPT) vs. ChatGPT – Non-switchers ($n=68$) Dependent Variable M (UniGPT) SD (UniGPT) M (ChatGPT) SD (ChatGPT) t df p Holm $\alpha$1 d2 95% CI3 Privacy concerns (H5) 2.32 0.89 3.46 0.80 -9.34 67 ¡ 0.001 0.00833 -1.13 [-1.44, -0.82] Experienced hallucinations (H4) 3.10 1.05 3.86 0.84 -6.41 67 ¡ 0.001 0.01 -0.78 [-1.05, -0.50] Trust (H1) 3.60 0.71 3.09 0.81 5.97 67 ¡ 0.001 0.0125 0.72 [0.45, 0.99] Sustainability-aware AI use (H6) 2.30 1.13 2.25 1.16 0.91 67 0.37 0.0167 0.11 [-0.13, 0.35] Cautious behavior toward hallucinations (H3) 4.20 0.67 4.19 0.66 0.10 67 0.92 n/a4 0.01 [-0.23, 0.25] Usefulness5 4.03 0.74 3.97 0.80 0.68 67 0.50 0.08 [-0.16, 0.32] Ease of use5 4.52 0.67 4.46 0.58 1.02 67 0.31 0.12 [-0.11, 0.36]
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below $\alpha$; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
Table D1. Mixed-Effects Model of system type, knowledge scores, and their interaction (number of observations: 232; $n=116$)
Effect Estimate SE t p
Trust: System $0.389^{**}$ $0.130$ $3.00$ $0.003$
Trust: System × Knowledge 0.049 $0.039$ $1.25$ $0.215$
CatiousBehavior: System -0.058 $0.084$ 0.69-0.69 $0.490$
CatiousBehavior: System × Knowledge 0.041 $0.025$ $1.61$ $0.111$
ExperiencedHallucinations: System $0.537^{***}$ $0.154$ 3.49-3.49 <0.001<0.001
ExperiencedHallucinations: System × Knowledge -0.027 $0.046$ 0.60-0.60 $0.553$
Privacy: System $1.228^{***}$ $0.193$ 6.38-6.38 <0.001<0.001
Privacy: System × Knowledge -0.001 $0.058$ 0.02-0.02 $0.987$
Sustainability: System 0.066 $0.101$ $0.65$ $0.514$
Sustainability: System × Knowledge 0.016 $0.030$ $0.53$ $0.594$
$^{*}p<.05$; $^{**}p<.01$; $^{***}p<.001$
All models include random intercepts for participants.
Appendix E Exploratory Analyses
Dependent Variable M (UniGPT) SD (UniGPT) M (ChatGPT) SD (ChatGPT) t df p Holm $\alpha$1 d2 95% CI3
Privacy concerns 2.47 0.96 3.17 0.90 -5.32 110.54 ¡ 0.001 0.0083 -0.76 [-1.03, -0.48]
Sustainability-aware AI use 2.88 1.21 2.19 1.03 4.28 103.37 ¡ 0.001 0.01 0.64 [0.36, 0.91]
Experienced hallucinations 3.40 1.05 3.75 0.94 -2.48 106.38 0.01 0.0125 -0.36 [-0.63, -0.09]
Cautious behavior toward hallucinations 4.29 0.66 4.10 0.74 1.95 129.60 0.05 0.0167 0.25 [-0.02, 0.52]
Trust 3.33 0.77 3.27 0.75 0.54 113.91 0.59 n/a4 0.07 [-0.19, 0.34]
Usefulness5 3.61 0.88 3.83 0.82 -1.84 109.78 0.07 -0.26 [-0.53, 0.01]
Ease of use5 4.46 0.66 4.60 0.59 -1.63 105.83 0.11 -0.24 [-0.51, 0.03]
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below α; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
[…]
Dependent Variable M (UniGPT) SD (UniGPT) M (ChatGPT) SD (ChatGPT) t df p Holm $\alpha$1 d2 95% CI3
Privacy concerns (H5) 2.47 0.96 3.17 0.93 -5.14 122.70 ¡ 0.001 0.0083 -0.74 [-1.02, -0.45]
Sustainability-aware AI use (H6) 2.88 1.21 2.19 1.05 4.19 112.56 ¡ 0.001 0.01 0.63 [0.35, 0.91]
Experienced hallucinations (H4) 3.40 1.05 3.77 0.98 -2.56 118.90 0.011 0.0125 -0.37 [-0.65, -0.09]
Cautious behavior toward hallucinations (H3) 4.29 0.66 4.15 0.71 1.39 136.89 0.17 0.0167 0.19 [-0.09, 0.47]
Trust (H1) 3.33 0.77 3.22 0.76 0.92 125.53 0.36 n/a4 0.13 [-0.15, 0.41]
Usefulness5 3.61 0.88 3.74 0.84 -1.04 121.41 0.30 -0.15 [-0.43, 0.13]
Ease of use5 4.46 0.66 4.59 0.59 -1.47 114.81 0.14 -0.22 [-0.50, 0.06]
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below α; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
Table E2. Welch’s independent t-test comparisons (between-subjects) of the customized LLMaaS chatbot (UniGPT) and ChatGPT Free, with participants using only one system ($n_{\text{UniGPT}}=70$, $n_{\text{ChatGPT Free}}=176$)